Trevor is interested in purchasing the local hardware/sporting goods store in the small town of Dove Creek, Montana. After ex(e) more than 17 out of the next 20 business days
Answer #1

-- - -- I Given Data: - 50 °၂ - 6.6 / P :-05 0.s ) D = ၆ , P( x ) - (x=3) +Ptx = 4) + F(x5) - ၀၈၀%”)+(၂၀၁၀+) - (o x8-125*62)+

d) n=20 ; fewer than fewer than 6 6 out out of the next of the next 20 business days p(X<6) – Pex-0) +P(X=1) +PQX=2) +P(x=3)